If 6b + 20 = a, and 4b + 30 = a, then b =
Answer by stanbon(75887)

If
6b + 20 = a
4b + 30 = a
--------
Substitute to solve for "b":
6b+20 = 5b+30
Subtract 5b from both sides to get:
b + 20 = 30
Subtract 20 from both sides to get:
b = 10
==================
Cheers,
Stan H.